OpenStudy (anonymous):

Which of the following has a slope of 2 and a y-intercept of –4?

OpenStudy (anonymous):

y = –2x + 4 y = 2x + 4 y = 2x – 4 y = –2x – 4

OpenStudy (jdoe0001):

\(\large \begin{array}{llll} y=&{\color{green}{ m}}&+{\color{blue}{ b}}\\ &\uparrow &\uparrow \\ &{\color{green}{ slope}}&{\color{blue}{ y-intercept}} \end{array}\) notice, slope = m = 2 y-intercept is -4

OpenStudy (anonymous):

thanks again

OpenStudy (jdoe0001):

yw